OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle?

OpenStudy (anonymous):

Have you learned Pythagorean's Theorem yet? It is \[a ^{2} + b ^{2} = c ^{2}\] where a, b, and c are the sides of a right triangle.

OpenStudy (anonymous):

if you kno so much why dont you give me the answer?

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (anonymous):

Just getting answers without knowing why is how homework can get done but it means you will do poorly on any exams if you don't know the methods.

OpenStudy (anonymous):

Not quite. Gaby got b^2. You need the square root of that.

OpenStudy (gabylovesyou):

yea what @mbradar2 said

OpenStudy (anonymous):

can i just get the answer with an explanation of how you got it ?

OpenStudy (theviper):

\[\Large{c^2=a^2+b^2}\]\[\Large{b^2=c^2-a^2}\]\[\Large{b=\sqrt{c^2-a^2}}\]Now put the values :)

OpenStudy (theviper):

\[\Large{b=\sqrt{12^2-11^2}}\]\[\Large{b=\sqrt{144-121}}\]\[\Large{b=\sqrt{23}}\]

OpenStudy (theviper):

Understood??

OpenStudy (theviper):

or\[b=4.80\space approximately\]